OpenStudy (miguelreyna19):

A combined circuit has two resistors in parallel (6.0 ohms and 8.0 ohms) and another in series (13.0 ohms). If the power source is 9.0 volts, what will the current be?

OpenStudy (jadzia):

You have to combine all the resistor first: For parallel resistors you use the formula: \(\sf 1/R_{eq} =1/R_1 + 1/R_2 + ...\) For series resistor you use the formula \(\sf R_{eq} = R_1 + R_2 + ...\) So for your problem, you have TWO parallel resistors, so calculate the equivalent resistor and then just add the other resistor.

OpenStudy (jadzia):

After you find the \(\sf R_{total}\) , you can apply Ohm's Law: \(\sf I= V/R_{total}\) , to get the current.